# Selection rule of canonical differential equations from Intersection theory

### v1

@@ Abstract
Canonical differential equations matrix is the 1-$\mathrm{d}\log$-form coefficients of projecting ($n$+1)-$\mathrm{d}\log$-forms onto $n$-$\mathrm{d}\log$-forms master integrand. Intersection number with dual form in relative cohomology could achieve the projecting and give selection rules of canonical differential equations from the pole structure of $\mathrm{d}\log$ master integrand.
@@ 1 Introduction
Integration-By-Parts (IBP) [1] and IBP-based differential equation [2, 3, 4, 5] method are two of the most powerful methods in perturbation quantum field theory. When we appropriately select the master integrals, differential equations of them could be transformed into canonical form [6],
@@
in which $\big{(}\mathrm{d}\Omega\big{)}_{IK}$ is $\mathrm{d}\log$-form proportional to $\epsilon$. Thereby it enables analytical solving master integrals as multi-polylogarithm [7, 8]. This method is called Canonical Differential Equations (CDE) and has become the most important analytic method in Feynman integrals’ evaluation. The $\mathrm{d}\log\text{W}^{(i)}(\bm{s})$ in (1) are referred to as symbol letters, whose complete set is known as symbol alphabet, constrains the function space of analytic solutions and can be utilized for bootstrap such as [9, 10, 11, 12, 13]. Thus, the properties of symbol letters or says also the $\mathrm{d}\log\text{W}^{(i)}(\bm{s})$ in (1) are important and have been studied in [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
One way to "appropriately select the master integrals" to get CDE is constructing $\mathrm{d}\log$ integrand as master integrands (which is also related to leading singularity) [6, 33, 34, 35, 36, 37, 38]. In [35, 33, 34], people have shown Baikov representation [39] make the construction more convenient. Recently, [40] shows how CDE emerges from $\mathrm{d}\log$ integrand of master integrals via intersection theory [41, 42]. In that work, only leading-order (LO) contribution and next-to-leading-order (NLO) contribution of intersection number are needed and both have universal formulas and selection rules of CDE have been given. However, in [40], the analyzing needs to keep some regulators for the computation of the intersection number, which should be taken zero at the end. The regulators lead to redundancy in both calculation and conclusion.
In this paper, supposing we have constructed $\mathrm{d}\log$ integrand as master integrals, we improve the selection rule of CDE in two aspects. Firstly, we show that CDE matrix is 1-$\mathrm{d}\log$-form coefficients of projecting ($n$+1)-$\mathrm{d}\log$-forms onto $n$-$\mathrm{d}\log$-forms master integrand, while such projecting only need use LO contribution of intersection number. Secondly, we apply the mathematical tool called twisted relative cohomology [43, 44, 45, 46] to avoid regulators, and redundancy as well, at the beginning. Thus we find the exact CDE selection rules.
@@ 2 Notations of Intersection Theory and Relative Cohomology
Intersection theory describe integrals take the form $I[u,\varphi]\equiv\int u\,\varphi\,,$ where
$\varphi\equiv\hat{\varphi}(\bm{z})\bigwedge_{j}\mathrm{d}z_{j}=\frac{Q(\bm{z})}{\left(\prod_{k}D_{k}^{a_{k}}\right)\left(\prod_{i}P_{i}^{b_{i}}\right)}\bigwedge_{j}\mathrm{d}z_{j}\,,\quad u=\prod_{i}\left[P_{i}(\bm{z})\right]^{\beta_{i}}\,,\quad a_{k},b_{j}\in\mathbb{N}\,,$ (2)
Feynman integrals and their representation, such as Baikov representation, Feynman parameterization, are integrals in this form.
Define the IBP-equivalence classes of cocycles [41, 42, 47, 48]:
$\bra{\varphi_{L}}\equiv\varphi_{L}\sim\varphi_{L}+\sum_{i}\nabla_{i}\xi_{i}\,,\,\quad\nabla_{i}=dz_{i}\wedge(\partial_{z_{i}}+\hat{\omega}_{i})\,,\quad\hat{\omega}_{i}\equiv\partial_{z_{i}}\log(u).$ (3)
which leads to a cohomology group. Any integral $f_{0}=\int u\varphi_{0}$ (or to be exact, the cocycle $\bra{\varphi_{0}}$ | corresponding to the integrals) in the function family are the vectors in space of cohomology group, and master integrals $f_{i}=\int u\varphi_{i}$ are the basis: $\bra{\varphi_{0}}=\sum_{i}c_{i}\bra{\varphi_{i}}.$ | . $c_{i}$ are the coefficients of reduction and given by projecting, instead of IBP
@@
where we use the IBP-invariant inner-product intersection number:
@@
In (5), $\bm{p}$ are the points (not lines or hypersurfaces with higher dimension) where n (or more in the case of degenerate pole) hypersurfaces belong to =$\mathcal{B}=\{P_{1}=0,\infty,\cdots,D_{1}=0,\infty,\cdots\}$ meet. ($P_{1}=0,\infty$ means $P_{1}=0,P_{1}=\infty$ ).
Notice that $D_{k}$ in (5) denote the denominators with integer power in $u\varphi$. In the case of Feynman integral, such denominators usually arise from propagators. With them, one could not directly compute (5). One way to handle them is adding regulators of them in u.
$u=\prod_{i}\left[P_{i}(\bm{z})\right]^{\beta_{i}}\prod_{k}D_{k}^{\delta_{k}}$ (6)
Then take the regulators to be zero in the final result. This was adopted in [40] and led to redundancy.
While in this paper, we will adopt the tool called relative cohomology to handle them. We do not go into mathematical details. Roughly speaking, we will let the dual forms $\varphi_{R}$ live on maximal cut of them, and denote them as $\Delta_{\hat{R}}\varphi_{R}$. In practice, it equivalent to applying the maximal cut of the sector of $\varphi_{R}$ to both sides before compute intersection number.
@@
Every symbol with the subscript $\hat{R}$ means it is cut version of the original one. $D_{R_{i}}$ in (7) usually are propagators of the sector of $\varphi_{R}$. If $\mathcal{B}_{\hat{I}}\subseteq\mathcal{B}_{\hat{J}}$, we say $\mathcal{B}_{\hat{I}}$ is a subsector of $\mathcal{B}_{\hat{J}}$.
To keep the information of multivariate pole explicitly in the calculation, we select to analyze and compute intersection number and the multivariate residue in it via multivariate Laurent expansion after factorization transformation corresponding to each poles [48, 40]
$\mathrm{T}^{(\alpha)}:z_{i}\to f^{(\alpha)}_{i}(\bm{x}^{(\alpha)})$
@@
Only after this step, one can apply multivariate expansion legally to $\varphi$ and solve intersection number
@@
where the powers $\bm{b}=(b_{1},\ldots,b_{n})$.
3 Selection Rule of CDE
@@ 3.1 $\mathrm{d}\log$ projection and LO intersection number
For distinction, we use $\mathrm{d}$ to denote the differentiation of integration variables $\bm{z}$ , and use $\hat{\mathrm{d}}$ to denote the differentiation of arbitrary one parameter, such as kinetic parameters or mass. And define $\mathrm{D}=\mathrm{d}+\hat{\mathrm{d}}$.
What we need is analyzing and computing CDE by (4) with (7)
@@
where $\varphi_{I}$ in original space and dual space are the same $\mathrm{d}\log$ integrand supposed to have been constructed $\varphi_{I}=\bigwedge_{j}\mathrm{d}\log\mathrm{W}^{(I)}_{j}(\bm{z})$. And $\dot{\varphi}_{I}$ in (10) is defined by $\hat{\mathrm{d}}f_{I}\equiv\int u\dot{\varphi}_{I}$. So notice that
@@
$\dot{\varphi}_{I}$ is ($n$+1)-$\mathrm{D}\log$-form. Then all we need for computing (10) is formula of LO contribution of intersection number (after factorization of poles)
$\operatorname{Res}_{\bm{\rho}^{(\alpha)}}\mathrm{T}^{(\alpha)}\left[\left(\nabla_{1}^{-1}\cdots\nabla_{n}^{-1}\varphi_{L}\right)\varphi_{R}\right]$ $=\frac{\operatorname{Res}_{\bm{\rho}^{(\alpha)}}\mathrm{T}^{(\alpha)}\left[\varphi_{L}\right]\times\operatorname{Res}_{\bm{\rho}^{(\alpha)}}\mathrm{T}^{(\alpha)}\left[\varphi_{R}\right]}{\prod_{i}\operatorname{Res}_{\rho^{(\alpha)}_{i}}\partial_{x_{i}^{(\alpha)}}\log\left(\mathrm{T}^{(\alpha)}\left[u\right]\right)\mathrm{d}x_{i}^{(\alpha)}}$
$=\frac{C_{L}^{(\bm{b}_{L})}C_{R}^{(\bm{b}_{R})}}{\bm{\gamma}^{(\alpha)}}\,,\quad\quad\bm{\gamma}^{(\alpha)}=\prod_{i}\gamma_{i}^{(\alpha)}$ (12)
@@ 3.2 Condition of $\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}\neq 0$
Take (12) and that $\varphi_{I}$ is $\mathrm{d}\log$-form into consideration, we could conclude the condition of $\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}\neq 0$
$\begin{cases}\mathrm{d}\log\text{-form}&:\ b_{I,i}\geq-1\,,\ b_{J,i}\geq-1\,,\\ \dot{\varphi}_{I;\hat{J}}&:\ b_{\dot{I},j}=b_{I,j}-1\,,\ \text{for one $j$}\,,\\ \braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}\neq 0&:\ b_{\dot{I},i}+b_{J,i}\leq-2\end{cases}$
$\Rightarrow\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}\neq 0:\ (-2\leq b_{I,j}+b_{J,j}\leq-1)\ \&\ (b_{I,i}=b_{J,i}=-1,i\neq j).$ (13)
We have omit the subscript "$;\hat{J}$" of $b_{*,*}$ for keeping the expression simple, but one should remember they all correspond to signs after maximal cut of sector $\mathcal{B}_{\hat{J}}$. We denote the two case as $\varphi_{I;\hat{J}}$ and $\varphi_{J;\hat{J}}$ share $n_{\hat{J}}$-variable Simple Pole ($n_{\hat{J}}$-SP): $(b_{I,j}+b_{J,j}=-1)\&(b_{I,i}=b_{J,i}=-1,i\neq j)$, and share ($n_{\hat{J}}$-1)-SP: $b_{I,i}=b_{J,i}=-1$.
For the case they share ($n_{\hat{J}}$-1)-SP, $\mathrm{D}\log\left(\mathrm{T}^{(\alpha)}[u]\right)$ in (11) provide pole of one remaining variable, and gives
$\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}=-\frac{\gamma^{(\alpha)}_{j;\hat{J}}}{\bm{\gamma}^{(\alpha)}_{\hat{J}}}\,\hat{\mathrm{d}}\int C_{I}^{(\bm{b}_{I})}C_{J}^{(\bm{b}_{J})}\,\hat{\mathrm{d}}\rho^{(\alpha)}_{j;{\hat{J}}}\,,$ (14)
For the case they share $n_{\hat{J}}$-SP, we have
$\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}=\frac{C_{I}^{(\bm{-1})}C_{J}^{(\bm{-1})}}{\bm{\gamma}^{(\alpha)}_{\hat{J}}}\,\hat{\mathrm{d}}\log\Big{(}\bar{u}_{\alpha;\hat{J}}(\bm{\rho}^{(\alpha)}_{\hat{J}})\Big{)}\,.$ (15)

### v2

# Selection rules of canonical differential equations from Intersection theory

@@ Abstract
The matrix of canonical differential equations consists of the 1-$\mathrm{d}\log$-form coefficients obtained by projecting ($n$+1)-$\mathrm{d}\log$-forms onto $n$-$\mathrm{d}\log$-form master integrands. With dual form in relative cohomology, the intersection number can be used to achieve the projection and provide the selection rules for canonical differential equations, which relate to the pole structure of the $\mathrm{d}\log$ master integrands.
@@ 1 Introduction
Integration-By-Parts (IBP) [1] and IBP-based differential equations [2, 3, 4, 5] are two of the most powerful methods in perturbative quantum field theory. By appropriately selecting the master integrals, their differential equations can be transformed into a canonical form [6]:
@@
In this expression, the $\big{(}\mathrm{d}\Omega\big{)}_{IK}$ is a $\mathrm{d}\log$-form proportional to $\epsilon$. This property allows for the analytical solution of master integrals as multi-polylogarithm [7, 8]. This method called Canonical Differential Equations (CDE) has become a crucial analytic tool for evaluating Feynman integrals. The $\mathrm{d}\log\text{W}^{(i)}(\bm{s})$ in (1) are referred to as symbol letters, and their complete set is known as the symbol alphabet. This set constrains the function space of analytic solutions and can be utilized for bootstrap techniques [9, 10, 11, 12, 13]. Therefore, the properties of symbol letters, or $\mathrm{d}\log\text{W}^{(i)}(\bm{s})$ in (1), are important and have been studied extensively in [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
One way to ’appropriately select the master integrals’ to obtain CDE is by constructing $\mathrm{d}\log$ integrand as master integrands. Such construction is also related to leading singularities [6, 33, 34, 35, 36, 37, 38]. In [35, 33, 34], it has been shown that Baikov representation [39] makes the construction of master integrands more convenient. Recently, [40] showed how CDE emerge from $\mathrm{d}\log$-form integrand of master integrals via intersection theory [41, 42]. In that work, one only needs to use leading-order (LO) contribution and next-to-leading-order (NLO) contribution of intersection number. These contributions have universal formulas, and based on them, selection rules for CDE have been provided. However, in [40], the analysis requires the use of regulators to compute the intersection numbers, which are supposed to be set to zero at the end. The inclusion of these regulators introduces redundancies in both the calculations and the conclusions.
In this paper, we enhance the selection rules for CDE presented in [40] in two aspects. Firstly, we present a new perspective on CDE by showing that the CDE matrix is the 1-$\mathrm{d}\log$-form coefficients obtained by projecting ($n$+1)-$\mathrm{d}\log$-forms onto $n$-$\mathrm{d}\log$-forms. This projection only requires the LO contribution to the intersection number, in contrast to the LO and NLO contributions used in [40]. Secondly, we use the mathematical tool of twisted relative cohomology [43, 44, 45, 46] to avoid regulators. This approach helps us to eliminate redundancy from the start. As a result, we obtain a more precise version of the CDE selection rules beyond [40].
@@ 2 Notations of Intersection Theory and Relative Cohomology
Intersection theory describes integrals of the form: $I[u,\varphi]\equiv\int u\,\varphi\,,$ where
$\varphi\equiv\hat{\varphi}(\bm{z})\bigwedge_{j}^{n}\mathrm{d}z_{j}=\frac{Q(\bm{z})}{\left(\prod_{k}D_{k}^{a_{k}}\right)\left(\prod_{i}P_{i}^{b_{i}}\right)}\bigwedge_{j}^{n}\mathrm{d}z_{j}\,,\quad u=\prod_{i}\left[P_{i}(\bm{z})\right]^{\beta_{i}}\,.$ (2)
Here $P_{i}$, $Q$ and $D_{k}$ are polynomials/monomials, $a_{k},b_{j}\in\mathbb{N}$ and $\beta_{i}\in\mathbb{C}$. $I[u,\varphi]$ encompasses various integral representations, including Baikov representation and Feynman parameterization of Feynman integrals.
Then, IBP relations can be rewrite as
$\int\sum_{i}\partial_{z_{i}}(u\xi_{i})\mathrm{d}z_{i}=\int u(\varphi+\sum_{i}\nabla_{i}\xi_{i})=0\,,\,\quad\nabla_{i}=dz_{i}\wedge(\partial_{z_{i}}+\hat{\omega}_{i})\,,\quad\hat{\omega}_{i}\equiv\partial_{z_{i}}\log(u)\,,$ (3)
where $u\xi_{i}$s are arbitrary integrands belong to the integral family but in $(n-1)$-form. Follow this expression, one can define the IBP-equivalence classes of cocycles [41, 42, 47, 48]:
$\bra{\varphi_{L}}\equiv\varphi_{L}\sim\varphi_{L}+\sum_{i}\nabla_{i}\xi_{i}\,.$ (4)
This leads to a cohomology group. Each integral $f_{0}=\int u\varphi_{0}$ corresponds to a vector in the cohomology group’s linear space [41, 42]. The master integrals $f_{i}=\int u\varphi_{i}$ form the basis of this space: $\bra{\varphi_{0}}=\sum_{i}c_{i}\bra{\varphi_{i}}.$ | . The coefficients $c_{i}$ are determined by projecting rather than using IBP:
@@
Here the $\braket{\varphi_{L}}{\varphi_{R}}$ denotes the intersection number, an IBP-invariant inner product. In (6), $\bm{p}$ represents the isolated points where $n$ hypersurfaces in =$\mathcal{B}=\{P_{1}=0,\infty,\cdots,D_{1}=0,\infty,\cdots\}$ intersect. Notice that $D_{k}$ in (6) represent the denominators with integer powers in $u\varphi$. In the context of Feynman integrals, such denominators typically arise from propagators. Direct computation of (6) can be challenging with these denominators. One approach to manage this issue is to introduce regulators $\delta_{k}$ for such denominators in u: $u=\prod_{i}\left[P_{i}(\bm{z})\right]^{\beta_{i}}\prod_{k}D_{k}^{\delta_{k}}$ Afterward, the regulators are set to zero to obtain the final result. This method was used in [40], but it led to redundancies.
In this paper, we use a technique known as relative cohomology to handle these denominators. While we do not delve into the mathematical details, the basic idea is as follows: we let the dual forms $\varphi_{R}$ live in the maximal cut of the denominators and denote them as $\Delta_{\hat{R}}\varphi_{R}$. Practically, this is equivalent to applying the maximal cut of the sector of $\varphi_{R}$ to both sides before computing the intersection number.
@@
Every symbol with the subscript $\hat{R}$ denotes the cut version of the original one. In (7), $D_{R_{i}}$ typically represent the propagators of the sector associated with $\varphi_{R}$. If $\mathcal{B}_{\hat{I}}\subseteq\mathcal{B}_{\hat{J}}$, we say that $\mathcal{B}_{\hat{I}}$ is a subsector of $\mathcal{B}_{\hat{J}}$.
To explicitly retain the information about multivariate poles in the calculations, we choose to analyze and compute the intersection number and the associated multivariate residue using a multivariate Laurent expansion. This expansion is applied after performing a factorization transformation corresponding to each pole [48, 40]:
$\mathrm{T}^{(\alpha)}:z_{i}\to f^{(\alpha)}_{i}(\bm{x}^{(\alpha)})\,,$
@@
Only after this step one can legally apply a multivariate expansion to $\varphi$ and compute the intersection number:
@@
Here, $\bm{b}=(b_{1},\ldots,b_{n})$ represents the powers in the Laurent expansion. When $b_{L,i}+b_{R,i}=-2$, we say it gives a LO contribution to intersection number. In such case, the contribution has a simple formula:
$\operatorname{Res}_{\bm{\rho}^{(\alpha)}}\mathrm{T}^{(\alpha)}\left[\left(\nabla_{1}^{-1}\cdots\nabla_{n}^{-1}\varphi^{(\bm{b}_{L})}_{L}\right)\varphi^{(\bm{b}_{R})}_{R}\right]=\frac{C_{L}^{(\bm{b}_{L})}C_{R}^{(\bm{b}_{R})}}{\tilde{\bm{\gamma}}^{(\alpha)}}\,,$
$\tilde{\bm{\gamma}}\equiv\prod_{i}\tilde{\gamma}_{i}\,,\quad\tilde{\gamma}_{i}^{(\alpha)}=\gamma_{i}^{(\alpha)}-b_{R,i}-1\,,\ \ \ \bm{b}_{L}+\bm{b}_{R}=\bm{-2}\,.$ (10)
3 Selection Rules of CDE
@@ 3.1 $\mathrm{d}\log$ projection and LO intersection number
For clarity, we use $\mathrm{d}$ to denote differentiation with respect to integration variables, and $\hat{\mathrm{d}}$ to denote differentiation with respect to one arbitrary parameter, such as kinetic parameters or masses. We define $\mathrm{D}=\mathrm{d}+\hat{\mathrm{d}}$.
To analyze and compute the CDE using (5) and (7), we start with the following equation:
@@
Here, $\varphi_{I}$ in both the original space and the dual space is the same $\mathrm{d}\log$ integrand, constructed as $\varphi_{I}=\bigwedge_{j}\mathrm{d}\log\mathrm{W}^{(I)}_{j}(\bm{z})$. The term $\dot{\varphi}_{I}$ in (11) is defined as $\hat{\mathrm{d}}f_{I}\equiv\int u\dot{\varphi}_{I}$. Therefore, we have:
@@
where $\dot{\varphi}_{I}$ is an ($n$+1)-$\mathrm{D}\log$-form. To compute (11), we only need to compute LO contributions of the intersection number (after factorizing the pole), which are given by (10).
@@ 3.2 Condition of $\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}\neq 0$
Considering (10) and that $\varphi_{I}$ is a $\mathrm{d}\log$-form, we can derive the condition under which $\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}\neq 0$:
$\begin{cases}\mathrm{d}\log\text{-form}&:\ b_{I,i}\geq-1\,,\ b_{J,i}\geq-1\\ \dot{\varphi}_{I;\hat{J}}&:\ b_{\dot{I},j}=b_{I,j}-1\\ \braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}\neq 0&:\ b_{\dot{I},i}+b_{J,i}\leq-2\end{cases}$
$\Rightarrow\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}\neq 0:\ (-2\leq b_{I,j}+b_{J,j}\leq-1)\ \&\ (b_{I,i}=b_{J,i}=-1,i\neq j)\,,$ (13)
for specific one $j$. Here we omit the subscript "$^{;\hat{J}}$" of $b_{*,*}$ to simplify the expression, but it should be noted that all $b_{*,*}$ here correspond to the powers after the maximal cut of the sector $\mathcal{B}_{\hat{J}}$. We denote two cases for $\varphi_{I;\hat{J}}$ and $\varphi_{J;\hat{J}}$ sharing poles as follows:
  • Sharing $n_{\hat{J}}$-Variable Simple Pole ($n_{\hat{J}}$-SP): This occurs when $b_{I,j}+b_{J,j}=-1$ for one specific $j$ and $b_{I,i}=b_{J,i}=-1$ for all $i\neq j$
  • Sharing $(n_{\hat{J}}-1)$-Variable Simple Pole ($(n_{\hat{J}}-1)$-SP): This occurs when $b_{I,i}=b_{J,i}=-1$.
For the case in which they share an ($n_{\hat{J}}$-1)-SP, $\mathrm{D}\log\left(\mathrm{T}^{(\alpha)}[u]\right)$ in (12) provides a pole of one remaining variable. The shared pole contributes
$-\frac{\gamma^{(\alpha)}_{j;\hat{J}}}{\bm{\gamma}^{(\alpha)}_{\hat{J}}}\,\hat{\mathrm{d}}\int C_{I}^{(\bm{b}_{I})}C_{J}^{(\bm{b}_{J})}\,\hat{\mathrm{d}}\rho^{(\alpha)}_{j;{\hat{J}}}$ (14)
to intersection number $\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}$. It is a $\hat{\mathrm{d}}\log$-form as discussed in [40].
For the case in which they share an $n_{\hat{J}}$-SP, the shared pole contributes
$\frac{C_{I}^{(\bm{-1})}C_{J}^{(\bm{-1})}}{\bm{\gamma}^{(\alpha)}_{\hat{J}}}\,\hat{\mathrm{d}}\log\Big{(}\bar{u}_{\alpha;\hat{J}}(\bm{\rho}^{(\alpha)}_{\hat{J}})\Big{)}$ (15)
to intersection number $\braket{\dot{\varphi}_{I;\hat{J}}}{\varphi_{J;\hat{J}}}$.